What is the value in year 13 of a $2,300 cash flow made in year 6 when the interest rates are 11 percent ?

Time= 13-6 = 7 years

Rate= 11%

PV= 1300

FV in 13 year= PV(1+R)t

FV= 2300(1.11)7= 4784

Is this correct? because this is not one of the answers
Answer #1

Answer : $4776.46

PV = Cash flow in Year 6 / FVIF(11%,6)
PV = 2300 / 1.8704 = $1230
Cash flow in Year 13 = PV * FVIF(11%,13)
Cash flow in Year 13 = 1230 * 3.8833 = $4776.46